OpenStudy (anonymous):

Find the limit. \[\lim_{h \rightarrow 0}\frac{ \frac{ 1 }{ (3+h)^2 }-\frac{ 1 }{ 9 } }{ h }\]

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

how far have you gotten in calculus?

OpenStudy (solomonzelman):

MMM... WHAT WAS YOUR FUNCTION TO BEGIN WITH ?

OpenStudy (anonymous):

Well, right now we're at extreme values, but we having a mid-term next week. So we're reviewing.

OpenStudy (solomonzelman):

k, now my question (lol)

OpenStudy (misty1212):

if you have gotten to derivatives, you recognize this as the derivative of \(f(x)=\frac{1}{x^2}\) at \(x=3\)

OpenStudy (misty1212):

take the derivative of \(f(x)=\frac{1}{x^2}\) then find \(f'(3)\) if you have not gotten to derivatives you have more work to do

OpenStudy (anonymous):

Can we not do this with derivatives, I have to review for limits. So the long way to solve this...

OpenStudy (anonymous):

@SolomonZelman I'm not sure what you mean. This is question, there is no other functions for this. That's the original...

OpenStudy (solomonzelman):

oh, that was the original question ?

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\lim_{h \rightarrow ~0}\frac{\frac{1}{(3+h)^2}-\frac{1}{9}}{h}}\) find the common denominator and add the fractions on the top.

OpenStudy (anonymous):

Yes, I did that

OpenStudy (solomonzelman):

and what do you get after doing this ?

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\ \frac{ \frac{ -6h-h^2 }{ 9(9+6h+h^2) } }{ h }\]

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\lim_{h \rightarrow ~0}\frac{\frac{1}{(3+h)^2}-\frac{1}{9}}{h}}\) \(\large\color{slate}{\displaystyle\lim_{h \rightarrow ~0}\frac{\frac{9}{9(3+h)^2}-\frac{(h+3)^2}{9(h+3)^2}}{h}}\) \(\large\color{slate}{\displaystyle\lim_{h \rightarrow ~0}\frac{\frac{9-(h+3)^2}{9(3+h)^2}}{h}}\) \(\large\color{slate}{\displaystyle\lim_{h \rightarrow ~0}\frac{9-(h+3)^2}{9h(3+h)^2}}\) \(\large\color{slate}{\displaystyle\lim_{h \rightarrow ~0}\frac{3^2-(h+3)^2}{9h(3+h)^2}}\) \(\large\color{slate}{\displaystyle\lim_{h \rightarrow ~0}\frac{(3-h-3)(3+h+3)}{9h(3+h)^2}}\)

OpenStudy (solomonzelman):

like this I am using a difference of squares (in my last step)

OpenStudy (solomonzelman):

(no need to expand the (3+h)^2 as you did, even though at times it appears to be an intuitive thing to do)

OpenStudy (solomonzelman):

if you still have questions, ask...

OpenStudy (anonymous):

okay... and then what? Sub in the 0, but it would be indeterminate...

OpenStudy (freckles):

3-h-3=?

OpenStudy (anonymous):

0?

OpenStudy (freckles):

3-h-3 3-3-h 0-h -h so 3-h-3=-h

OpenStudy (freckles):

\[\large\color{slate}{\displaystyle\lim_{h \rightarrow ~0}\frac{(3-h-3)(3+h+3)}{9h(3+h)^2}} \\ \large\color{slate}{\displaystyle\lim_{h \rightarrow ~0}\frac{(-h)(3+h+3)}{9h(3+h)^2}} \\ \]

OpenStudy (freckles):

you should see something that cancels

OpenStudy (anonymous):

OH! I see! the h's reduce and it would be -6/81 --> -2/27

OpenStudy (freckles):

yah

OpenStudy (solomonzelman):

yes, it is correct. freckles tnx for continuing for me.

OpenStudy (anonymous):

Thank you! @SolomonZelman and @freckles

OpenStudy (solomonzelman):

yw